# DiffLM: Controllable Synthetic Data Generation via Diffusion Language Models

### v1

@@ 4.2 Code Generation
Benchmarking. In the code generation scenario, to simplify the problem, we focus on Python code and use the Flytech [^1] https://huggingface.co/datasets/flytech/python-codes-25k dataset as real data, which contains 24,813 unique real user queries and the corresponding Python code fulfilling those requests. We discard the user queries and use only the code to train DiffLM. After generating synthetic code data, we continue pre-training the Mistral 7B v0.3 base model (Jiang et al., 2023) using a smaller learning rate, i.e., 1e-5, in a causal language modeling objective. We then benchmark the trained model on code generation tasks, selecting two mainstream benchmarks: HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). To better understand the performance changes of the base model, we also experiment with base models of different sizes, i.e., Mistral Nemo with 12B parameters.
Baselines. We include baselines from recent code models. First, we consider the CodeLLaMA (Rozière et al., 2023) series, which use approximately 600B tokens to continue pre-training the LLaMA-2 (Touvron et al., 2023) base model, injecting strong code capabilities through multi-task learning. Additionally, we compare with the Mistral base model (Jiang et al., 2023) and its instruction-tuned variants, the latter could representing the upper bound of code capabilities for this architecture.
Evaluation. We report the code generation capabilities in Table 2. Specifically, Mistral-Real-Code and Mistral-DiffLM-Code denote models that were further pre-trained on real data and synthetic data generated by DiffLM, respectively. The 7B models are based on Mistral-0.3-Base, and the 12B models are based on Mistral-Nemo-Base. Both models were trained for 3 epochs on the same amount of data using identical hyperparameters, effectively serving as a controlled experiment where the data source is the only variable. The results indicate that simply continuing to pre-train the Mistral model with a small amount of code data leads to inconsistent impacts on code generation capabilities. Specifically, Mistral-Real-Code shows a slight improvement on HumanEval but a significant decline on MBPP. However, using our synthetic data to continue pre-training the base model yields better results than using real data. For instance, Mistral-DiffLM-Code-7B, achieved a 7 percentage point improvement over the base model, even outperforming the Code Llama 7B model that was trained with more extensive data. In summary, in the code generation scenario, we focus on the differing impacts of real data and synthetic data, further demonstrating that DiffLM can generate synthetic data that is even more effective than real data in enhancing downstream task performance.
@@ 4.3 Tool Generation
Evaluation. To address more complex structured data generation scenarios, we further conduct a tool synthesis task. Specifically, we select the ToolBench (Qin et al., 2024) dataset as a benchmark for comparison, which is constructed based on the RapidAPI [^2] https://rapidapi.com/hub platform by crawling APIs created by real users and synthesizing related dialogue SFT data using GPT [^3] https://chat.openai.com. We use the its toolset to train DiffLM and then sample an equal number of tools for comparison. We assess the usability of the generated tools from two perspectives: 1) Single-Tool Quality: We use GPT-4 as an annotator to score the real and synthetic data across multiple dimensions on a scale from 0 to 10, where the results are illustrated in Figure 2. 2) Category-Level Preference: We collect all tools within the same category and use GPT-4 to perform preference scoring between real tools and synthetic tools, as presented in Table 3. The specific evaluation prompts are provided in the appendix B.2. From the results, DiffLM’s synthetic data achieves higher scores in the single-tool scoring task, indicating that leveraging the internal knowledge and generative capabilities of LLMs allows us to create tool descriptions and input/output parameter definitions of higher textual quality. Additionally, in the category-level preference evaluation, nearly $1/3$ of the tool types surpass or are on par with real data in terms of diversity and usability. Since DiffLM can sample and generate tools indefinitely to increase coverage, we believe there is room for further improvement in this metric.
@@ 5 Analysis
Figure 3: Model loss curves under different latent feature injection methods and different $\beta$ adjustment strategies. The left is the KL-divergence loss trends, and the right is the language modeling reconstruction loss on a logarithmic scale. In the cyclical $\beta$ strategy, $\beta$ increases linearly from 0 to 0.2. The other methods employ a decreasing $\beta$, starting from a maximum value of 0.1 and decreasing to a minimum of 0.001. Our proposed injection and $\beta$ adjustment strategy achieves the lowest reconstruction loss.
5.1 Ablation Study
@@ 6 Conclusion
In this paper, we introduce DiffLM, a novel framework designed to enhance LLM’s understanding of real-world data distributions in synthetic data generation tasks. DiffLM leverages a VAE to map real data into a latent space, which is then injected into the decoding process of LLM, enabling end-to-end training through causal language modeling objective. Additionally, we incorporate a diffusion process to further refine the learning of the latent distribution, mitigating the sampling failures caused by latent space discrepancies. To flexibly and non-intrusively control the structure and quality of the generated data, DiffLM integrates real data information with LLMs’ internal knowledge by freezing the LLM parameters and using the latent features as plug-in modules. Experimental results demonstrate that DiffLM produces highly robust and consistent outputs. In all datasets, the performance of downstream models trained on the generated data is comparable to or even surpasses that of models trained on real data.
@@ Appendix A Details on Model Design
A.1 Diffusion Process
@@
The reverse diffusion process aims to recover $\mathbf{z}_{0}$ from a noisy latent variable $\mathbf{z}_{t}\sim\mathcal{N}(0,\mathbf{I})$. It is parameterized by a neural network $\bm{\epsilon}_{\theta}$, which predicts the noise component at each timestep:
@@
Typically, the model predicts the mean $\mu_{\theta}$ while the covariance $\Sigma_{\theta}$ is fixed or simplified. By leveraging the properties of the forward process, the mean can be parameterized to predict the original noise $\bm{\epsilon}$ added during the forward diffusion:
$\mu_{\theta}(\mathbf{z}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)\right).$ (13)
@@
The training objective for LDMs focuses on minimizing the difference between the true noise $\bm{\epsilon}$ added during the forward process and the noise predicted by the model $\bm{\epsilon}_{\theta}$. The simplified loss function is:
$\mathcal{L}_{\text{latent}}=\mathbb{E}_{\mathbf{x}_{0},\bm{\epsilon},t}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)\right\|^{2}\right],$ (14)
@@
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\,\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}).$ (15)
@@ B.3 Instructions for Reproduction
Appendix C Synthetic Data Generated by DiffLM
In Table 6, we compare the real test data of the Adult dataset with the generated outputs from GReaT and DiffLM. As discussed in Section 5.1, DiffLM produces more diverse samples that more closely align with the real data distribution. Specifically, for columns like workclass and native-country, the outputs generated by the GReaT model are relatively homogeneous.

### v2

@@ 4.2 Code Generation
Benchmarking. In the code generation scenario, to simplify the problem, we focus on Python code and use the Flytech [^1] https://huggingface.co/datasets/flytech/python-codes-25k dataset as real data, which contains 24,813 unique real user queries and the corresponding Python code fulfilling those requests. We discard the user queries and use only the code to train DiffLM. After generating synthetic code data, we continue pre-training the Mistral 7B v0.3 base model (Jiang et al., 2023) using a smaller learning rate, i.e., 1e-5, in a causal language modeling objective. We then benchmark the trained model on code generation tasks, selecting two mainstream benchmarks: HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021b). To better understand the performance changes of the base model, we also experiment larger models like Mistral Nemo with 12B parameters.
Baselines. We include the performance from recent code LLMs. First, we consider the CodeLLaMA (Rozière et al., 2023) series, which use approximately 600B tokens to continue pre-training the LLaMA-2 (Touvron et al., 2023) base model, injecting strong code capabilities through multi-task learning. Additionally, we compare with the Mistral base model (Jiang et al., 2023) and its instruction-tuned variants, the latter could representing the upper bound of code capabilities for this architecture.
Evaluation. We report the code generation capabilities in Table 2. Specifically, Mistral-Real-Code and Mistral-DiffLM-Code denote models that were further pre-trained on real data and synthetic data generated by DiffLM, respectively. The 7B models are based on Mistral-0.3-Base, and the 12B models are based on Mistral-Nemo-Base. Both models were trained for 3 epochs on the same amount of data using identical hyperparameters, effectively serving as a controlled experiment where the data source is the only variable. The results indicate that simply continuing to pre-train the Mistral model with a small amount of code data leads to inconsistent impacts on code generation capabilities. Specifically, Mistral-Real-Code shows a slight improvement on HumanEval but a significant decline on MBPP. However, using our synthetic data to continue pre-training the base model yields better results than using real data. For instance, Mistral-DiffLM-Code-7B, achieved a 7% improvement over the base model, even outperforming the Code Llama 7B model that was trained with more extensive data. In summary, in the code generation scenario, we focus on the differing impacts of real data and synthetic data, further demonstrating that DiffLM can generate synthetic data that is even more effective than real data in enhancing downstream task performance.
@@
Figure 3: Model loss curves under different latent feature injection methods and different $\beta$ adjustment strategies. The left is the KL-divergence loss trends, and the right is the language modeling reconstruction loss on a logarithmic scale. In the cyclical $\beta$ strategy, $\beta$ increases linearly from 0 to 0.2. The other methods employ a decreasing $\beta$, starting from a maximum value of 0.1 and decreasing to a minimum of 0.001.
@@ 4.3 Tool Generation
Evaluation. To address more complex structured data generation scenarios, we further conduct a tool synthesis task. Specifically, we select the ToolBench (Qin et al., 2024) dataset as a benchmark for comparison, which is constructed based on the RapidAPI [^2] https://rapidapi.com/hub platform by crawling APIs created by real users and synthesizing related dialogue SFT data using GPT [^3] https://chat.openai.com. We use the its toolset to train DiffLM and then sample an equal number of tools for comparison. We assess the usability of the generated tools from two perspectives: 1) Single-Tool Quality: We use GPT-4 as an annotator to score the real and synthetic data across multiple dimensions on a scale from 0 to 10, where the results are illustrated in Figure 2. 2) Category-Level Preference: We collect all tools within the same category and use GPT-4 to perform preference scoring between real tools and synthetic tools, as presented in Table 4.2. The specific evaluation prompts are provided in the appendix B.2. From the results, DiffLM’s synthetic data achieves higher scores in the single-tool scoring task, indicating that leveraging the internal knowledge and generative capabilities of LLMs allows us to create tool descriptions and input/output parameter definitions of higher textual quality. Additionally, in the category-level preference evaluation, nearly $1/3$ of the tool types surpass or are on par with real data in terms of diversity and usability. Since DiffLM can sample and generate tools indefinitely to increase coverage, we believe there is room for further improvement in this metric.
@@ 6 Conclusion
In this paper, we propose DiffLM, a novel framework that enhances LLMs’ ability to generate high-quality synthetic data by integrating real-world data distributions. DiffLM maps real data into a latent space via VAE, which is then injected into LLM decoding through a causal language modeling objective. A diffusion process further refines the latent distribution, reducing sampling discrepancies. To ensure flexible and non-intrusive control over data quality and structure, DiffLM freezes LLM parameters and employs latent features as plug-in modules. Experiments show that DiffLM generates consistent data, enabling downstream models to match or even exceed the performance of those trained on real data. For future research, our work could help build data flywheels for LLM applications and facilitate dataset curation for low-resource domains.
Limitations
Our work, while offering valuable contributions to synthetic data generation, has several limitations. First, our approach focuses primarily on structured data synthesis, such as tabular data generation, code generation, and tool generation. It does not address the generation of more subjective or nuanced data, such as sentiment-specific data or tasks requiring varied levels of complexity or capability. Additionally, our decoupling of data distribution learning may need further refinement to accommodate such subjective text synthesis tasks. Second, due to time and hardware constraints, we limited our study to a relatively simple VAE and diffusion network, and used LLMs up to 7B parameters for decoding. Future research could explore the potential of scaling these models further, using even larger architectures to better evaluate and validate the effectiveness of the DiffLM framework in data synthesis tasks.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62272439) and the Fundamental Research Funds for the Central Universities, and is conducted during Ying Zhou’s internship at ByteDance Inc.
@@ Appendix A Details on Model Design
A.1 Controllability
We define the controllability of text data synthesis as the ability to generate text that satisfies desired requirements (e.g., structure, topics, domains) (Keskar et al., 2019; Li et al., 2022). Existing methods for structured textual data synthesis often struggle with controllability. On one hand, LLM prompt-based methods relying on prompt engineering or few-shot inference cannot guarantee the diversity and scalability of synthetic data, even with complex human-crafted processes (Long et al., 2024). On the other hand, controlling a LM by fine-tuning it with supervised data (SFT, RLHF) is not only expensive but might also degrade the LLM’s general capability (Keskar et al., 2019; Borisov et al., 2023). Our method addresses these challenges through sampling in the latent space while maintaining data structure due to LLM’s instruction-following ability.
A.2 Diffusion Process
@@
The reverse diffusion process aims to recover $\mathbf{z}_{0}$ from a noisy latent variable $\mathbf{z}_{t}\sim\mathcal{N}(0,\mathbf{I})$. It is parameterized by a neural network $\boldsymbol{\epsilon}_{\theta}$, which predicts the noise component at each timestep:
@@
Typically, the model predicts the mean $\mu_{\theta}$ while the covariance $\Sigma_{\theta}$ is fixed or simplified. By leveraging the properties of the forward process, the mean can be parameterized to predict the original noise $\boldsymbol{\epsilon}$ added during the forward diffusion:
$\mu_{\theta}(\mathbf{z}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t)\right).$ (12)
@@
The training objective for LDMs focuses on minimizing the difference between the true noise $\boldsymbol{\epsilon}$ added during the forward process and the noise predicted by the model $\boldsymbol{\epsilon}_{\theta}$. The simplified loss function is:
$\mathcal{L}_{\text{latent}}=\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon},t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t)\right\|^{2}\right],$ (13)
@@
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\,\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}).$ (14)
@@ Appendix B Details on Experimental Setup
B.1 Tabular Data Generation
Table 5 shows the detail information of tabular dataset we use in paper.
@@ B.3 Instructions for Reproduction
B.4 Training Parameters for Baselines.
We reproduced the tabular results using the code released by the original paper, ensuring that all hyperparameters and settings were consistent with the original implementation. All results were almost identical to those reported in the TabSyn paper; therefore, we used the results reported in TabSyn in Table 1 to ensure a fair comparison.
Appendix C Addition Experiment Results
C.1 Training Data Plagiarism
C.2 Visualization
Table 6: Tabular MLE performance with different VAE decoder models for DiffLM.
Method Adult Default Magic Shoppers Beijing
Real 0.927 - 0.770 - 0.946 - 0.926 - 0.423 -
GReaT 0.913 12.12 0.755 19.94 0.888 16.16 0.902 14.51 0.653 8.25
ICL-Lllama 0.848 16.02 - - 0.792 21.6 0.871 27.56 1.005 17.71
DiffLM-Llama 0.879 8.89 0.785 9.67 0.914 8.41 0.906 6.93 0.707 10.11
ICL-Mistral 0.803 29.22 0.732 36.00 0.881 25.64 0.882 42.00 0.865 12.45
DiffLM-Mistral 0.906 9.74 0.794 9.06 0.917 7.53 0.915 10.07 0.696 6.35
C.3 DiffLM with other VAE Decoders
In Table 6, we report the tabular MLE performance of DiffLM using LLaMA3-8B as the VAE decoder. As illustrated in the table, DiffLM-LLaMA consistently outperforms the ICL-based method across all metrics and datasets. Specifically, on the Default dataset, while the ICL method was unable to generate valid synthetic samples, DiffLM successfully synthesized high-quality data, achieving an MLE of 0.785. This result surpasses the MLE performance of real data (0.770), clearly indicating DiffLM’s capability to produce synthetic data of superior quality compared to authentic samples. On the Shoppers dataset, DiffLM-LLaMA achieves an MLE of 0.906, closely approaching the 0.915 obtained by DiffLM-Mistral. Furthermore, DiffLM-LLaMA attains a rho value of 6.93, outperforming DiffLM-Mistral’s rho value of 10.07. This demonstrates that DiffLM can deliver even more favorable outcomes when integrated with more advanced language models.
Table 7: Tabular MLE performance with varying quantity of real and synthetic data. Performance on the Beijing dataset is evaluated using the RMSE metric, where lower values indicate better performance. 2x means we use double training synthesized data for evaluation.
Adult Default Magic Shoppers Beijing
Real 0.927 0.770 0.946 0.926 0.423
TabSyn (SoTA) 0.915 0.764 0.938 0.920 0.582
DiffLM (1x) 0.894 0.793 0.910 0.9122 0.717
DiffLM (2x) 0.896 0.795 0.914 0.9124 0.704
(+0.002) (+0.002) (+0.004) (+0.0002) (-0.013)
Real+DiffLM 0.925 0.802 0.936 0.932 0.494
C.4 Quantity of Synthetic Data.
We experimented with increasing the amount of data synthesized by DiffLM and combining real data with DiffLM-synthesized data. As shown in Table 7, adding more synthesized data further improves around 0.2% MLE performance in the tabular scenario. Since our method can synthesize unlimited amounts of data and we did not design any complex post-processing method, the performance improvement brought by DiffLM-synthesized data in downstream tasks still has significant room for growth. Additionally, combining real and synthetic data generated by DiffLM can improve downstream performance; all results exceed $>0.2$% of those using only DiffLM data. Notably, on the Beijing and Shoppers datasets, the combination of real data and DiffLM synthetic data surpasses 0.6%-3% of the performance of training on real data alone.
Table 8: The human evaluation results on 100 pairs of randomly selected DiffLM-generated tool and real tool within the same category. Averaged by 3 human experts with computer science knowledge.
Percentage
DiffLM Win 88%
Equal 6%
Real Win 6%
C.5 Human Evaluation.
We have used GPT-4 to rate and perform preference judgments on synthesized tools and real tools. The results in Figure 2 and Table 4.2 demonstrate the quality of our synthesized data. As per your suggestion, we have conducted human evaluations on the tools data. Specifically, we compared 100 pairs of randomly selected DiffLM-generated data and real data within the same category. As shown in Table 8, our synthetic data is preferred by human annotators.
C.6 More Result of Baselines
GReaT.
We attempted to validate the GReaT method on Mistral but found it could not directly and effectively generate data with the desired structure. GReaT organizes tabular data in a “key is value" format and uses a smaller PLM (i.e., GPT-2) for continued pretraining. However, when applied to larger models like Mistral, GReaT struggled to effectively generate the desired structured data. The sample generated by GReaT with Mistral is shown in Figure 9. We hypothesize that controlling an LM by fine-tuning it with supervised data cause catastrophic forgetting for LLMs, as suggested by Luo et al. (2023b). Specifically, the “key is value" data constructed by the GReaT method, when used to continue pre-training Mistral, causes internal knowledge collapse - both undermining the model’s existing knowledge and failing to do effective data synthesis. Additionally, training the adult dataset on GReaT for 200 epochs (default settings) requires approximately 50 hours on 8 A100 80G GPUs, which is resource-intensive. In contrast, DiffLM under the same training settings requires only about 7 hours.
TabSyn.
We want to emphasize that our goal is a unified structured data synthesis framework that supports various domains like tabular data, codes, and tools, and tabular data generation in our work is just a subdomain of synthetic data generation. As a comparison, TabSyn is not applicable to more complex data synthesis tasks, such as code generation and real-world tool generation, which involve generating longer content, more complex data types, and highly structured data, while our DiffLM can handle complicated scenarios. The results of tabular data synthesis are to demonstrate that our method possesses generality and can achieve on-par results with domain-specific models without being specifically tailored to a particular domain.
@@ 32 Local-gov Bachelors Adm-clerical Asian-Pac-Islander Male India $>50K$
C.7 Synthetic Data Generated by DiffLM
In Table 9, we show the samples from the Adult dataset with the generated tabular rows from GReaT and DiffLM method. As discussed in Section 5, DiffLM produces more diverse samples that more closely align with the real data distribution. Specifically, for columns like workclass and native-country, the outputs generated by the GReaT model are relatively homogeneous.
<s> income is <=50K, \$300/month income is >50K, \$600/month I am trying to figure out how to implement the above logic with the following table I have: | ID | Amt | |----|-----| | 1 | 450 | | 2 | 650 | | 3 | 500 | | 4 | 150 | | 5 | 700 | | 6 | 400 | I have tried using IF ELSE function in Excel but I am not sure how to apply it to this particular situation. You can use the IF function in Excel to implement the logic you provided. Here is a formula that you can use: ‘‘‘ =IF(A2<=50000, 300, 600) ‘‘‘ In this formula, A2 is the cell where you have the income. You can adjust the formula to apply it to your table by using the following formula: ‘‘‘ =IF(A2<=50000, 300, 600) ‘‘‘ Assuming that the income values are in column A, you can use the following formula to get the monthly payment for each ID: ‘‘‘ =IF(A2<=50000, 300, 600) ‘‘‘ Drag this formula down to get the payment for each ID. Here is an example of how your table would look like after applying the formula: | ID | Amt | Payment | |----|-----|---------| | 1 | 450 | 300 | | 2 | 650 | 600 | | 3 | 500 | 300 | | 4 | 150 | 300 | | 5 | 700 | 600 | | 6 | 400 | 300 |</s></s></s></s></s></s></s></s></s></s></s></s></s></s></s></s></s></s></s></s></s></s></s></s></s>’, ’<s> income is <=50K, credit score is >=700, and the applicant has lived in their current residence for more than 2 years. To create a segment based on the provided conditions in a SQL-like syntax, you can use the following: ‘‘‘sql SELECT * FROM customers WHERE income <= 50000 AND credit_score >= 700 AND length_of_residence > 2; ‘‘‘ This SQL query selects all records from the customers table where the income is less than or equal to 50,000, the credit score is 700 or higher, and the length of residence (assuming that length_of_residence is a field indicating the number of years a customer has lived at their current address) is more than 2.</s>
Figure 9: A random synthetic sample generated by GReaT trained with Mistral. Use the exactly same training and generating settings as GReaT with trained with GPT-2.